John applied to an accounting firm and a consulting firm. He knows that 30% of similarly qualified applicants receive job offers from the accounting firm, while only 20% of similarly qualified applicants receive job offers from the consulting firm. Assume that receiving an offer from one firm is independent of receiving an offer from the other. What is the probability that both firms offer John a job?
Answer #1

Answer)

Lets say we have a sample size of 100

30% of similarly qualified applicants receive job offers from the accounting firm, = 30

20% of similarly qualified applicants receive job offers from the consulting firm = 20

Probability = favorable/total

Total = 100

Favorable = 30+20 = 50

Required probability is 50/100 = 0.5